# Multilinear Wavelet Compact T(1) Theorem

### v2

@@ 4. Compact wavelet forms
for some $\sigma\in\mathrm{S}_{m}$ and $\nu_{z}\in\Psi^{m,\delta;j}_{z}$; by $\nu_{z}(x,\cdot)$ we mean for each $x$ it returns the function $(x_{2},\ldots,x_{m})\mapsto\nu_{z}(x,x_{2},\ldots,x_{m})$. We will only handle the second case, and we will reduce, by symmetry, to the case where $\sigma$ is the identity. The first case in (4.1) is simpler, though in fact they are handled in exactly the same fashion. Let $\rho>0$ and split the integral defining $T$ over $\mathbb{B}_{M}$ and $\mathbb{Z}^{d}\setminus\mathbb{B}_{M}$ where $M$ is chosen large enough that $|\varepsilon_{z}|\leq\rho$ for $z\not\in\mathbb{B}_{M}$. Therefore, the operator norm of the second component is, by (2.9) controlled by $\rho$. The proof will be concluded if we can show $R_{\rho}$ defined by
@@
is compact. By the Riesz-Kolmogorov compactness criteria (see e.g. [MSWW] and [Tsuji] for the case $10<r<$ 1), we need to prove that
@@
whence $\{T[\mathbf{f}_{n_{k}}]\}_{k\in\mathbb{N}}$ is Cauchy and has a limit in $\mathrm{CMO}(\mathbb{R}^{d})$. Now it remains to show each $R_{\rho}$ is compact. Applying (4.5) with all $p_{j}=\infty$ implies the same pointwise estimates as above, which implies $R_{\rho}:{\mathcal{B}}^{\infty}\to{\mathrm{C}}_{\mathrm{v}}(\mathbb{R}^{d})$ and that the family $\{R_{\rho}\mathbf{f}:\mathbf{f}\in{\mathcal{B}}^{\infty}\}$ is equicontinuous. Therefore, by the Arzela-Ascoli theorem and a diagonalization argument, given a sequence $\{\mathbf{f}_{n}\}_{n\in\mathbb{N}}\in{\mathcal{B}}^{\infty}$, we can obtain a subsequence such that $\{R_{\rho}\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ is Cauchy in $\left\|\cdot\right\|_{L^{\infty}([-n,n]^{d})}$ for each $n\in\mathbb{N}$. However, the pointwise estimate for $R_{\rho}\mathbf{f}$ shows that given $\epsilon>0$ we can find $n$ large enough that $\left|R_{\rho}\mathbf{f}_{n_{k}}(x)\right|<\frac{\epsilon}{3}$ for $x$ outside $[-n,n]^{d}$. Combining these two facts with the triangle inequality shows that $\{R_{\rho}\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ is Cauchy in $\left\|\cdot\right\|_{L^{\infty}}$ which is a stronger norm than $\left\|\cdot\right\|_{\mathrm{BMO}(\mathbb{R}^{d})}$. Finally, recalling that $R_{\rho}\mathbf{f}_{n_{k}}\in{\mathrm{C}}_{\mathrm{v}}(\mathbb{R}^{d})$ establishes that the limit belongs to $\mathrm{CMO}(\mathbb{R}^{d})$. ∎
@@ 5. Compactness of paraproducts
and the natural modification at the endpoint which we denote by $\left\|\cdot\right\|_{\mathrm{ess}(\infty,\mathrm{BMO})}$ where the operator norm of $T-K$ is measured in $\left\|\cdot\right\|_{\infty,\mathrm{BMO}}$.
@@
Since $\psi_{I}\in\Psi^{1,1;1}_{\zeta(I)}$ and $\phi_{z}\in\Psi^{1,1;1}_{z}$, one can compute $\left|\langle\psi_{I},\phi_{z}\rangle\right|\lesssim[z,\zeta(I)]_{\frac{1}{2}}$ (see e.g. [DWW]*Lemma 2.3 or [FJW]*Appendix, Lemmata 2 and 4). Furthermore, since $\zeta(I)\in\mathbb{B}_{M}$ and $z\not\in\mathbb{B}_{100M}$, $[z,\zeta]_{\frac{1}{2}}\lesssim M^{-\frac{1}{4}}[z,\zeta]_{\frac{1}{4}}$. Now, to apply Lemma 2.5, rewrite
@@ 6. Main result and proof
It is trivial that E. implies A., and the path (A. $\implies$ B. $\implies$ C.) is precisely the content of Proposition 3.8. To show C. implies D. we isolate the cancellative part of $\Lambda$, namely
@@
    For any $\varphi\in L^{\infty}_{c}$,
@@
Only A. and C. differ from their multilinear statements in Theorem 2. To compare C. with its $m$-linear analogue, one only needs to note that $z\in\mathbb{B}_{M}(\zeta)$ is equivalent to $\zeta\in\mathbb{B}_{M^{\prime}}(z)$ as long as $M\sim M^{\prime}$. Therefore, (B. $\implies$ C. $\implies$ D. $\implies$ E.) follows from Theorem 2. E. trivially implies A., so it remains to establish A. implies B. where the true force of the linear situation is used. Namely, $t^{\frac{d}{2}}\phi_{z}$ converges weakly to zero and therefore, since $T$ is compact $TT^{*}(t^{\frac{d}{2}}\phi_{z})$ and $T^{*}T(t^{\frac{d}{2}}\phi_{z})$ both converge strongly to zero, which implies B. ∎

### v3

@@ 4. Compact wavelet forms
for some $\sigma\in\mathrm{S}_{m}$ and $\nu_{z}\in\Psi^{m,\delta;j}_{z}$; by $\nu_{z}(x,\cdot)$ we mean for each $x$ it returns the function $(x_{2},\ldots,x_{m})\mapsto\nu_{z}(x,x_{2},\ldots,x_{m})$. We will only handle the second case, and we will reduce, by symmetry, to the case where $\sigma$ is the identity. The first case in (4.1) is simpler, though in fact they are handled in exactly the same fashion. Let $\rho>0$ and split the integral defining $T$ over $\mathbb{B}_{M}$ and $Z^{d}\setminus\mathbb{B}_{M}$ where $M$ is chosen large enough that $|\varepsilon_{z}|\leq\rho$ for $z\not\in\mathbb{B}_{M}$. Therefore, the operator norm of the second component is, by (2.10) controlled by $\rho$. The proof will be concluded if we can show $R_{\rho}$ defined by
@@
is compact. By the Riesz-Kolmogorov compactness criteria (see e.g. [MSWW] and [Tsuji] for the case $0<r<1$), we need to prove that
@@
whence $\{T[\mathbf{f}_{n_{k}}]\}_{k\in\mathbb{N}}$ is Cauchy and has a limit in $\mathrm{CMO}(\mathbb{R}^{d})$. Now it remains to show each $R_{\rho}$ is compact. Applying (4.5) with all $p_{j}=\infty$ implies the same pointwise estimates as above, which implies $R_{\rho}:{\mathcal{B}}^{\infty}\to{\mathrm{C}}_{\mathrm{v}}(\mathbb{R}^{d})$ and that the family $\{R_{\rho}\mathbf{f}:\mathbf{f}\in{\mathcal{B}}^{\infty}\}$ is equicontinuous. Therefore, by the Arzela-Ascoli theorem and a diagonalization argument, given a sequence $\{\mathbf{f}_{n}\}_{n\in\mathbb{N}}\in{\mathcal{B}}^{\infty}$, we can obtain a subsequence such that $\{R_{\rho}\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ is Cauchy in $\left\|\cdot\right\|_{L^{\infty}([-n,n]^{d})}$ for each $n\in\mathbb{N}$. However, the pointwise estimate for $R_{\rho}\mathbf{f}$ shows that given $\epsilon>0$ we can find $n$ large enough that $\left|R_{\rho}\mathbf{f}_{n_{k}}(x)\right|<\frac{\epsilon}{3}$ for $x$ outside $n,n]^{d}$. Combining these two facts with the triangle inequality shows that $\{R_{\rho}\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ is Cauchy in $\left\|\cdot\right\|_{L^{\infty}}$ which is a stronger norm than $\left\|\cdot\right\|_{\mathrm{BMO}(\mathbb{R}^{d})}$. Finally, recalling that $R_{\rho}\mathbf{f}_{n_{k}}\in{\mathrm{C}}_{\mathrm{v}}(\mathbb{R}^{d})$ establishes that the limit belongs to $\mathrm{CMO}(\mathbb{R}^{d})$. ∎
@@ 5. Compactness of paraproducts
and the natural modification at the endpoint which we denote by $\left\|\cdot\right\|_{\mathrm{ess}(\infty,\mathrm{BMO})}$ where the operator norm of TKT-K is measured in $\left\|\cdot\right\|_{\infty,\mathrm{BMO}}$.
@@
Since $\psi_{I}\in\Psi^{1,1;1}_{\zeta(I)}$ and $\phi_{z}\in\Psi^{1,1;1}_{z}$, one can compute $\left|\langle\psi_{I},\phi_{z}\rangle\right|\lesssim[z,\zeta(I)]_{\frac{1}{2}}$ (see e.g. [DWW]*Lemma 2.3 or [FJW]*Appendix, Lemmata 2 and 4). Furthermore, since $\zeta(I)\in\mathbb{B}_{M}$ and $z\not\in\mathbb{B}_{100M}$, $z,\zeta]_{\frac{1}{2}}\lesssim M^{-\frac{1}{4}}[z,\zeta]_{\frac{1}{4}}$. Now, to apply Lemma 2.5, rewrite
@@ 6. Main result and proof
It is trivial that E. implies A., and the implication from A. to B. and B. to C. is precisely the content of Proposition 3.8. To show C. implies D. we isolate the cancellative part of $\Lambda$, namely
@@
    For any $\varphi\in L^{\infty}$ with compact support,
@@
Only A. and C. differ from their multilinear statements in Theorem 2. To compare C. with its $m$-linear analogue, one only needs to note that $z\in\mathbb{B}_{M}(\zeta)$ is equivalent to $\zeta\in\mathbb{B}_{M^{\prime}}(z)$ as long as $M\sim M^{\prime}$. Therefore, the implications from B. to C., from C. to D., and from D. to E. all follow from Theorem 2. E. trivially implies A., so it remains to establish A. implies B. where the true force of the linear situation is used. Namely, $t^{\frac{d}{2}}\phi_{z}$ converges weakly to zero and therefore, since $T$ is compact $TT^{*}(t^{\frac{d}{2}}\phi_{z})$ and $T^{*}T(t^{\frac{d}{2}}\phi_{z})$ both converge strongly to zero, which implies B. ∎